What is the boiling point of benzene under certain pressure condition is 80.1°c .calculate the boiling point of a solution containing 5g of solute x in 100g of benzene under the same pressure condition given that KB = 2.6°c per 100g of benzene and molar mass of solute x is 229.
1
Expert's answer
2020-11-09T13:50:09-0500

Boiling point of solution T(solution) is:

T(solution) = T(benzene) + Kb*molarity

molarity = 5 g / 229 g/mole / 0.100 kg = 0.218 mole/kg

T(solution) = 80.1 + 0.218*2.6 = 80.7 °c

Boiling point of the solution is 80.7 °c